# Rotatable Array-Aided Hybrid Beamforming for Integrated Sensing and Communication

### v1

@@ Abstract
Integrated Sensing and Communication (ISAC) is one of the pivotal supporting technologies for next-generation wireless communication networks. As an emerging technical means, the new six-dimensional movable antenna (6DMA) system can effectively improve the communication and sensing performance of ISAC systems. However, related research on the architecture based on three-dimensional rotatable antennas (RA) remains relatively limited. Especially under the influence of non-ideal channel characteristics, how to balance system performance and hardware cost control through beamforming technology in non-ideal channels has become a crucial issue to be solved urgently. Given the significant advantages of hybrid beamforming technology in balancing system performance and hardware complexity, this paper focuses on the channel model considering effective aperture loss under the RA architecture and studies the sub-connected hybrid beamforming design for multi-user ISAC systems. Aiming at the non-convex nature with coupled variables in this problem, this paper first transforms the complex fractional objective function using the Fractional Programming (FP) framework, and then proposes an algorithm based on the Alternating Optimization (AO) architecture, which achieves optimization by alternately solving five subproblems. Among them, the closed-form update expression of the array rotation angle is derived through the Karush-Kuhn-Tucker (KKT) conditions, and a two-stage Gradient Ascent (GA) based method is proposed to optimize the antenna rotation angle. Simulation results show that compared with the traditional fixed-position antenna (FPA), the proposed method can significantly improve the overall performance of the system.
@@ I Introduction
Integrated sensing and communication (ISAC) has become a highly promising technology in the sixth-generation (6G) wireless network [1][2]. Its core advantage is the ability to share hardware platforms and signal bandwidths. This helps achieve efficient resource use for communication and radar sensing. This feature not only greatly improves spectrum efficiency, but also cuts hardware costs a lot [3]. Because of these advantages, ISAC has shown great potential in many applications. These include vehicle-to-everything (V2X), industrial internet of things (IIoT), and environmental monitoring [2]. In addition, ISAC has attracted wide attention from both academia and industry. It integrates a series of emerging wireless technologies, such as reconfigurable intelligent surface (RIS) [4] [5], unmanned aerial vehicle (UAV) [6], and non-orthogonal multiple access (NOMA) [7][8]. This further enhances its performance and application potential.
@@
With the continuous evolution of 6G communication systems, the large-scale application of high-frequency millimeter wave and large antenna array has become an irreversible technological trend. By increased antenna numbers, large antenna array can significantly improve beamforming gain and enhance spatial resolution, thereby effectively improving the communication and sensing performance. However, the fully-digital beamforming scheme requires each antenna unit to be equipped with an independent radio frequency chain, which makes the deployment of large antenna array face the challenge of a sharp rise in hardware costs and power consumption, seriously restricting its popularization and application in practical scenarios. In contrast, the hybrid beamforming [9][10], relying on a small number of RF chains and an analog front end consisting of phase shifters (PSs), has been considered as a promising technique to achieve a good trade-off between cost and communication performance.
@@
Based on the front end design of PS circuits, the hybrid beamformer can be categorized into fully-connected structures and sub-connected structures [11]. In the fully-connected structure, all RF chains are connected to all antenna elements through PSs, which endows a higher system degree of design freedom (DoFs). However, it also results in high hardware costs and energy consumption overheads, which are particularly prominent in the scenario of large-scale antenna array in millimeter wave systems. In contrast, the sub-connected structure, by connecting each RF chain only to a part of the antenna elements, has higher energy efficiency and is easier for engineering implementation. To improve system performance under the constraint of a limited number of RF chains, numerical optimization algorithms have been proposed[9][7][13]. [9] designs a fully-digital beamformer, then iteratively optimizes the hybrid beamformer to gradually approximate the performance of the fully-digital solution. [7] proposes a method to directly optimize the hybrid beamformer in an iterative manner. [13] introduces a continuous auxiliary variable to replace the beamforming variable with unit modulus constraint, optimizes the auxiliary variable, extracts its phase, and finally obtains the design result of the beamformer. Nevertheless, most of the existing fully-connected and sub-connected hybrid beamforming schemes rely on fixed position antennas (FPA), and the transmitter/receiver fail to fully exploit the spatial variation of wireless channels, thus ignoring the gain improvement that can be brought about by dynamic channel adjustment.
@@
ISAC aims primarily to enhance communication capacity and sensing ability [14]. In recent years, movable antenna (MA) and six-dimensional MA (6DMA) technologies have been proposed to improve the performance of wireless communication [20]-[19]. Specifically, compared with the traditional MA which can only dynamically adjust the position of antenna elements within a specific feasible region, the 6DMA can flexibly regulate its spatial position and rotation angle, and integrate the DoFs of the antenna in three-dimensional (3D) position and rotation dimensions, thus providing customized sensing and communication services more efficiently [16]-[19]. Several studies have investigated the integration of MA to enhance communication capacity [20]-[27]. Specifically, the authors analyze the performance of the MA enabled wireless communication system in [20]. They derive closed-form expressions for channel gain under deterministic and random channel environments, verifying the effectiveness of MA in improving channel capacity. [21] investigates a MA-enabled full-duplex ISAC system, which considers the scenario of user uplink and downlink transmissions and sets the sum of user uplink/downlink communication rates and target sensing rate as the optimization objective. The authors in [22] conduct research on ISAC systems with clutter interference, deriving in detail the derivative of channel gain with respect to antenna position, and optimizing the antenna position based on the gradient ascent method. The authors in [23][24] propose a scheme for MA aided multi-user communication scenarios under hybrid precoding design. The authors in [25] study the trade-off between communication and sensing in MA-enabled ISAC system, and adopts the Particle Swarm Optimization (PSO) algorithm to determine the optimal position of the antenna. [26] explores the application of MA in multi-target detection, and the results show that MA can significantly reduce transmission power while ensuring communication and sensing performance. [27] investigates multi-user communication systems supported by the collaboration of RIS and MA.
@@
Although the advantages of MA in ISAC systems have been validated, effectively enhancing communication and sensing performance by optimizing antenna positions remains a challenging issue. The main difficulty arises from the complex expression of the wireless channel with respect to the positions of the antennas, causing non-convex expressions involving coupling variables. Existing studies mostly focus on MA-enabled ISAC systems with adjustable positions. In contrast, the investigation of ISAC systems considering antenna rotatability is still in its infancy. Generally, antenna rotation can optimize beam pointing more flexibly, and can provide customized sensing and communication services more efficiently, especially when considering an accurate channel model that includes effective aperture [28]. To this end, this paper studies the hybrid beamforming design for multi-user multi-input-single-output (MU-MISO) ISAC systems based on rotatable antennas, aiming to reduce hardware cost and complexity while improving the communication and sensing performance of the system. To the best of the authors’ knowledge, this is among the first works to explore RA-based ISAC.The main contributions of this paper are summarized as follows.
@@
    We study a hybrid beamforming design for a RA enabled ISAC system. In this design, to balance sensing accuracy and communication efficiency, the goal is to maximize the weighted sum of each user’s communiations rate and the target sensing rate by jointly optimizing the transmit digital beamformer, transmit analog beamformer, receive beamformer, and antenna rotation angles.
@@
    We propose an alternating optimization (AO) based framework to solve the formulated problem, which enables the direct design of the hybrid beamformer. Specifically, the minimum mean square error (MMSE) filtering method is first adopted to optimize the receive beamformer. Then, Facilitated by fractional programming (FP), the remaining problem is decomposed into four subproblems to handle the coupled variables. Among them, an algorithm based on singular value decomposition (SVD) and Karush-Kuhn-Tucker (KKT) conditions is proposed to solve the analog beamforming problem, and the gradient ascent (GA) method is used to solve the array rotation angle optimization problem.
@@
    We perform numerical simulations to evaluate the performance of RA in ISAC system with our proposed algorithm, and demonstrate the performance trade-off between communication and sensing. Simulation results confirm that adopting RA with our proposed algorithm significantly enhances both sensing and communication performance with our proposed algorithm. Moreover, under certain practical conditions, the proposed scheme with sub-connected RA yields a higher sum rate compared to the fully-connected FPA array.
@@ II System Model
As illustrated in Fig. 1, we consider an RA-enabled far-field MU-MISO system. The system comprises a dual functional radar and communication (DFRC) BS serving $K$ single-FPA users, sensing one target, and considering $C$ clutters as interference for sensing. We consider a bistatic setup at the BS, which sparated Tx and Rx to avoid challenging issues in full-duplex mode. Thus, self-interference (SI) can be significantly suppressed through the Tx and Rx are sufficiently separated. Therefore, in this paper, we assume that the SI is completely cancellation. To reduce power consumption and hardware complexity of the BS, the BS employs the sub-connected hybrid array with $N_{t}$ transmit antennas, which is divided into $N_{RF}$ nonoverlapped subarrays, each containing $M=N_{\mathrm{t}}/N_{RF}$ antennas connected to an RF chain through analog phase shifters. We assume $M$ is an integer for brevity.
@@
II-A RA-BS Model
In this paper, we consider an RA-enabled far-field ISAC system. The BS is equipped with $N_{t}$ transmit antennas and $N_{r}$ receive antennas, which are respectively fixed on a rotatable transmit antenna plate and a rotatable receive antenna plate, each with an area of $D$. To avoid coupling effects, we ensure that the distance between any two antennas is always no less than half a wavelength. Without loss of generality, we assume that the antennas lie within the yzy-z plane of the local Cartesian coordinate system defined by the center of the corresponding plane, and the local positions of the $N_{t}$-th transmit antenna and the $N_{r}$-th receive antenna can be respectively represented as
@@ II-A RA-BS Model
$\mathbf{q}_{n_{r}}^{\mathrm{r}}=[0,y^{\mathrm{r}}_{n_{r}},z^{\mathrm{r}}_{n_{r}}]^{\mathrm{T}}\in A_{r},$ (2)
The transmit and receive array are connected to the CPU via rotatable rods embedded with flexible wires, and thus 3D rotations can be adjusted. The 3D rotation can be respectively characterized by
$\bm{\varpi}^{\mathrm{q}}=[\alpha^{\mathrm{q}},\beta^{\mathrm{q}},\gamma^{\mathrm{q}}]^{\mathrm{T}},q\in\{\mathrm{t},\mathrm{r}\}$ (3)
@@
where $\mathbf{C}^{\mathrm{t}}$ and $\mathbf{C}^{\mathrm{r}}$ denote the centers’positions of transmit plate and receive plate, respectively.
@@ II-B Channel Model
Consider a far-field, quasi-static channels, we assume the communication user (CU) and the target are located in the far-field region of the BS. Since the size of each UPA’s region is much smaller than the signal propagation distance between the BS and the scatter points and/or receiver, the planar wave is adopted to construct the field-response array model. The pointing vector of the CU/target corresponding to the direction $(\theta,\phi)$ are given by
@@
where $\theta\in[0,\pi]$ and $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$ denote the elevation and azimuth angles, respectively. The difference in signal path distance between the target position $\mathbf{p}$ to the $n_{t}$-th transmit antenna and from $\mathbf{p}$ to the $n_{r}$-th receive antenna is represented by $\mathbf{p}^{T}\mathbf{t}_{n_{\mathrm{t}}}(\bm{\varpi}^{t})$ and $\mathbf{p}^{T}\mathbf{r}_{n_{\mathrm{r}}}(\bm{\varpi}^{r})$, and the corresponding complex form of phase difference are given by $e^{-j\frac{2\pi}{\lambda}\mathbf{p}^{T}\mathbf{t}_{n_{\mathrm{t}}}(\bm{\varpi}^{t})}$ and $e^{-j\frac{2\pi}{\lambda}\mathbf{p}^{T}\mathbf{r}_{n_{\mathrm{r}}}(\bm{\varpi}^{r})}$. Taking $\mathbf{p}$ as the reference point, the steer vector of the Tx is
@@
As mentioned in [19], the rotatability of both transmit and receive array must meet some rotation constraints to avoid mutual signal reflections between them. But in this paper, we consider adopting a bistatic setup at the BS, where the Tx and the Rx are separated. Meanwhile, we keep a relatively large distance between the Tx and the Rx to avoid this problem. Therefore, we assume that there is no problem of mutual reflection of signals.
@@
Consequently, The communication channel vector $\mathbf{h}_{k}\in\mathbb{C}^{N_{t}\times 1}$ between the $k$-th CU and the Tx can be derived as
@@
and also, the field response vector from the BS to the sensing target/clutters can be written as $\mathbf{G}_{\chi}\in\mathbb{C}^{N_{r}\times N_{t}}$ between the BS and the target is modeled as
@@
where $\beta_{\chi}$ represents the target radar cross section (RCS) with $\mathbb{E}\{|\beta_{\chi}|^{2}\}=\zeta^{2}$, $\chi=\{s,1,\dots,C\}$ representing the target or the clutter, $\mathbf{g}_{\chi}$ and $\mathbf{h}_{\chi}$ represent the channel vector between the transmitter and target and between the receiver and target, respectively.

### v2

@@ Abstract
Six-dimensional movable antenna (6DMA) technology has been proposed to enhance the performance of Integrated Sensing and Communication (ISAC) systems. However, within 6DMA-related research, studies on the ISAC system based on rotatable array (RA) remains relatively limited. Given the significant advantages of hybrid beamforming technology in balancing system performance and hardware complexity, this paper focuses on a channel model that accounts for the efficiency of the antenna radiation pattern and studies the sub-connected hybrid beamforming design for multi-user RA-aided ISAC systems. Aiming at the non-convex nature with coupled variables in this problem, this paper transforms the complex fractional objective function using the Fractional Programming (FP) method, and then proposes an algorithm based on the Alternating Optimization (AO) framework, which achieves optimization by alternately solving five subproblems. For the analog beamforming optimization subproblem, we adopt Singular Value Decomposition (SVD) method to transform the objective function, thereby deriving the closed-form update expression for the analog beamforming matrix. For the antenna rotation optimization subproblem, we derive the closed-form derivative expression of the array rotation angle and propose a two-stage Gradient Ascent (GA) based method to optimize the antenna rotation angle. Extensive simulation results demonstrate the effectiveness of the proposed RA-aided hybrid beamforming design method. It not only significantly improves the overall system performance while reducing hardware costs, but also achieves performance comparable to that of the fully-digital beamforming design with fixed-position antennas (FPA) under specific parameter configurations.
@@ I Introduction
Integrated sensing and communication (ISAC) has become a highly promising technology in the sixth-generation (6G) wireless network [1][2]. Its core advantage is the ability to share hardware platforms and signal bandwidths, which helps achieve efficient resource use for communication and radar sensing. This feature not only greatly improves spectrum efficiency, but also cuts hardware costs substantially [3]. Due to these advantages, ISAC has shown great potential in many applications. These include vehicle-to-everything (V2X), industrial internet of things (IIoT), and environmental monitoring [2]. In addition, ISAC has attracted wide attention from both academia and industry. It integrates a series of emerging wireless technologies, such as reconfigurable intelligent surface (RIS) [4] [5], unmanned aerial vehicle (UAV) [6], and non-orthogonal multiple access (NOMA) [7][8]. These technologies further enhance ISAC’s performance and application potential.
@@
With the continuous evolution of 6G communication systems, the large-scale application of high-frequency millimeter wave and large antenna array has become an irreversible technological trend. By increasing antenna numbers, large antenna array can significantly improve beamforming gain and enhance spatial resolution, thereby effectively improving the communication and sensing performance. However, the fully-digital beamforming scheme requires each antenna unit to be equipped with an independent radio frequency chain, which makes the deployment of large antenna array face the challenge of a sharp rise in hardware costs and power consumption, seriously restricting its popularization and application in practical scenarios. In contrast, the hybrid beamforming [9][10], relying on a small number of RF chains and an analog front-end consisting of phase shifters (PSs), has been considered as a promising technique to achieve a good trade-off between cost and communication performance.
@@
Based on the front-end design of PS circuits, the hybrid beamformer can be categorized into fully-connected structures and sub-connected structures [11]. In the fully-connected structure, all RF chains are connected to all antenna elements through PSs, which endows a higher system degree of design freedom (DoFs). However, it also results in high hardware costs and energy consumption overheads, which is particularly prominent in the scenario of large-scale antenna array in millimeter wave systems. In contrast, the sub-connected structure, by connecting each RF chain only to a part of the antenna elements, has higher energy efficiency and is easier for engineering implementation. To improve system performance under the constraint of a limited number of RF chains, numerical optimization algorithms have been proposed[7][9][13]. [7] proposes a method to directly optimize the hybrid beamformer in an iterative manner. [9] designs a fully-digital beamformer, then iteratively optimizes the hybrid beamformer to gradually approximate the performance of the fully-digital solution. [13] introduces a continuous auxiliary variable to replace the beamforming variable with a unit modulus constraint, optimizes the auxiliary variable, extracts its phase, and finally obtains the design result of the beamformer. Nevertheless, most of the existing fully-connected and sub-connected hybrid beamforming schemes rely on fixed position antennas (FPA), and the transmitter/receiver fails to fully exploit the spatial variation of wireless channels, thus ignoring the gain improvement that can be brought about by dynamic channel adjustment.
@@
ISAC aims primarily to enhance communication capacity and sensing ability [14]. In recent years, movable antenna (MA) and six-dimensional MA (6DMA) technologies have been proposed to improve the performance of wireless communication [15]-[19]. Specifically, compared with the traditional MA which can only dynamically adjust the position of antenna elements within a specific feasible region, the 6DMA can flexibly regulate its spatial position and rotation angle, and integrate the DoFs of the antenna in three-dimensional (3D) position and rotation dimensions, thus providing customized sensing and communication services more efficiently [16]-[19]. Several studies have investigated the integration of MA to enhance communication capacity [20]-[27]. Specifically, the authors analyze the performance of the MA-enabled wireless communication system in [20]. They derive closed-form expressions for channel gain under deterministic and random channel environments, verifying the effectiveness of MA in improving channel capacity. [21] investigates an MA-enabled full-duplex ISAC system, which considers the scenario of user uplink and downlink transmissions and sets the sum of user uplink/downlink communication rate and target sensing rate as the optimization objective. The authors in [22] conduct research on ISAC systems with clutter interference, deriving in detail the derivative of channel gain with respect to antenna position, and optimizing the antenna position based on the gradient ascent method. The authors in [23][24] propose a scheme for MA-aided multi-user communication scenarios under hybrid precoding design. The authors in [25] study the trade-off between communication and sensing in MA-enabled ISAC system, and adopts the Particle Swarm Optimization (PSO) algorithm to determine the optimal position of the antenna. [26] explores the application of MA in multi-target detection, and the results show that MA can significantly reduce transmission power while ensuring communication and sensing performance. [27] investigates multi-user communication systems supported by the collaboration of RIS and MA.
@@
Although the advantages of MA in ISAC systems have been validated, effectively enhancing communication and sensing performance by optimizing antenna positions remains a challenging issue. The main difficulty arises from the complex expression of the wireless channel with respect to the positions of the antennas, causing non-convex expressions involving coupling variables. Existing studies mostly focus on MA-enabled ISAC systems with adjustable positions. In contrast, the investigation of ISAC systems considering antenna rotatability is still in its infancy. Generally, antenna rotation can optimize beam pointing more flexibly, and can provide customized sensing and communication services more efficiently, especially when considering a channel model that accounts for the efficiency of the antenna radiation pattern [28]. Thus, this paper studies the hybrid beamforming design for multi-user multi-input-single-output (MU-MISO) ISAC systems based on rotatable arrays, aiming to reduce hardware cost and complexity while improving the communication and sensing performance of the system. To the best of the authors’ knowledge, this is among the first works to explore RA-based ISAC systems. The main contributions of this paper are summarized as follows.
@@
    We study a hybrid beamforming design for an RA-aided ISAC system, where the channel model accounts for the efficiency of the antenna radiation pattern. To fully exploit the performance gains brought by the array rotation to the ISAC system, flexibly rotatable transmit and receive antenna panels are deployed at the BS to provide communication services for multiple users while sensing a target. We aim to maximize the weighted sum of each user’s communications rate and the target sensing rate by jointly optimizing the transmit digital beamformer, transmit analog beamformer, receive beamformer, and antenna rotation angles to balance sensing accuracy and communication efficiency.
@@
    We propose an alternating optimization (AO) method to solve the formulated non-convex optimization problem with highly-coupled variables. Specifically, the minimum mean square error (MMSE) filtering method is first adopted to optimize the receive beamformer. Then, facilitated by fractional programming (FP), the remaining problem is decomposed into four subproblems to handle the coupled variables. For the hybrid beamforming optimization, we adopt singular value decomposition (SVD) and Karush-Kuhn-Tucker (KKT) conditions to derive the closed-form solutions, thereby solving the analog beamforming optimization problem and the digital beamforming optimization problem. Then the closed-form derivative expression of the array rotation angle is derived, and a gradient ascent (GA) method is proposed to solve the array rotation angle optimization problem.
@@
    We perform numerical simulations to evaluate the advantages of sub-connected RA-aided ISAC system and the effectiveness of the proposed algorithm. Simulation results demonstrate that the RA-aided ISAC system outperforms the FPA-based ISAC system due to the additional DoFs introduced by array rotation optimization. Moreover, under certain practical conditions, the sub-connected RA-aided ISAC system yields higher communication and sensing performance compared to the fully-connected FPA-based system and the fully-digital FPA-based system.
@@ II System Model
II-A RA-BS Model
As illustrated in Fig. 1, this paper considers an RA-aided far-field MU-MISO system. The system comprises a dual-functional radar and communication (DFRC) BS serving $K$ single-FPA users, sensing one target, and considering $C$ clutters as interference for sensing. The BS is equipped with $N_{t}$ transmit antennas and $N_{r}$ receive antennas, which are respectively fixed on a rotatable transmit antenna panel (TP) and a rotatable receive antenna panel (RP), each with an area of $D$. To reduce power consumption and hardware complexity of the BS, the BS employs the sub-connected hybrid array, which is divided into $N_{RF}$ nonoverlapped subarrays, each containing $M=N_{\mathrm{t}}/N_{RF}$ antennas connected to an RF chain through PSs. $M$ is an integer for brevity. A bistatic setup is adopted at the BS, where TP and RP are separated to avoid challenging issues in full-duplex mode. Thus, self-interference (SI) can be significantly suppressed when the TP and RP are sufficiently separated. Therefore, this paper assumes that the SI is completely cancelled.
@@ II-A RA-BS Model
To avoid coupling effects, the distance between any two antennas is always no less than half a wavelength. Without loss of generality, the antennas lie within the yzy-z plane of the local Cartesian coordinate system defined by the center of the corresponding plane, and the local positions of the $N_{t}$-th transmit antenna and the $N_{r}$-th receive antenna can be respectively represented as
@@
$\mathbf{q}_{n_{r}}^{\mathrm{r}}=[0,y^{\mathrm{r}}_{n_{r}},z^{\mathrm{r}}_{n_{r}}]^{\mathrm{T}}\in A_{r}.$ (2)
The transmit and receive arrays are connected to the CPU via rotatable rods embedded with flexible wires, and thus 3D rotations can be adjusted. The 3D rotation can be respectively characterized by
$\bm{\varpi}^{\mathrm{q}}=[\alpha^{\mathrm{q}},\beta^{\mathrm{q}},\gamma^{\mathrm{q}}]^{\mathrm{T}},q\in\{\mathrm{t},\mathrm{r}\},$ (3)
@@
where $\mathbf{C}^{\mathrm{t}}$ and $\mathbf{C}^{\mathrm{r}}$ denote the centers’positions of TP and RP, respectively.
@@ II-B Channel Model
Considering a far-field quasi-static channel, the communication users (CUs) and the target are located in the far-field region of the BS. Since the size of each UPA’s region is much smaller than the signal propagation distance from the BS to the scatter points/receivers, the planar wave is adopted to construct the field-response array model. The pointing vector of the CUs/target corresponding to the direction $(\theta,\phi)$ is given by
@@
where $\theta\in[0,\pi]$ and $\phi\in[-\frac{\pi}{2},\frac{\pi}{2}]$ denote the elevation and azimuth angles, respectively. The difference in signal path distances from the target position $\mathbf{p}$ to the $n_{t}$-th transmit antenna and from $\mathbf{p}$ to the $n_{r}$-th receive antenna are represented by $\mathbf{p}^{T}\mathbf{t}_{n_{\mathrm{t}}}(\bm{\varpi}^{t})$ and $\mathbf{p}^{T}\mathbf{r}_{n_{\mathrm{r}}}(\bm{\varpi}^{r})$, and the corresponding complex forms of phase difference are given by $e^{-j\frac{2\pi}{\lambda}\mathbf{p}^{T}\mathbf{t}_{n_{\mathrm{t}}}(\bm{\varpi}^{t})}$ and $e^{-j\frac{2\pi}{\lambda}\mathbf{p}^{T}\mathbf{r}_{n_{\mathrm{r}}}(\bm{\varpi}^{r})}$. Taking $\mathbf{p}$ as the reference point, the steer vector of the TP is
@@
As mentioned in [19], the rotatability of both transmit and receive arrays must meet certain rotation constraints to avoid mutual signal reflections between them. In this paper, however, a relatively large distance is maintained between the TP and the RP to address this issue. Therefore, it is assumed that there is no mutual signal reflection.
@@
Consequently, The communication channel vector $\mathbf{h}_{k}\in\mathbb{C}^{N_{t}\times 1}$ between the $k$-th CU and the TP can be derived as
@@
and also, the field response vector from the BS to the sensing target/clutters can be written as $\mathbf{G}_{\chi}\in\mathbb{C}^{N_{r}\times N_{t}}$, which is modeled as
@@
where $\beta_{\chi}$ represents the target radar cross section (RCS) with $\mathbb{E}\{|\beta_{\chi}|^{2}\}=\zeta^{2}$, $\chi=\{s,1,\dots,C\}$ represents the target or the clutters, $\mathbf{g}_{\chi}$ and $\mathbf{h}_{\chi}$ represent the channel vector between the transmitter and target as well as between the receiver and target, respectively.